1. A) Suppose 30 % of home trade is with country 1 and 70 % is with country 2; Home’s currency appreciates 15 % against country 1 but depreciates 20 % against country 2. What is the change of effective exchange rate for the home country? Show your work.

B) What do you understand the concept of ”Absolute Purchasing Power Parity”? What does it mean when ”it holds”? Write it on your own understanding. Then, show the formula for ”Absolute Purchasing Power Parity” and explain it by using a numerical example.

C) Think of Asset Approach and consider we have an equilibrium in Foreign Exchange Rate Market (or Forex). Then consider the following situation: Home Interest Rate (Policy Rate) decreases. How does (spot) exchange rate change? Show your work on a graph and explain the intuition step by step.
Answer #1

Given

Home Trade
Country 1 - 30%
Country 2- 70%

Home currency appreciates
Country 1 - 15%
Country 2 - 20%

Change of Effective Exchange rate for the home country
1.   The real effective exchange rate (REER) is the weighted average of a country's currency in relation to an index or basket of other major currencies.

2.   Weights are determined by comparing the relative trade balance of a country's currency against each country.

3.   This exchange rate is used to determine an individual country's currency value relative to the other major currencies in the index.


Real Effective Exchange Rate (REER)
The Formula for REER Is

REER=CERn×CERn×CERn×100
where: CER = Country exchange rate
REER= 0.30*0.70*100
REER=21


2. Absolute Purchasing Power Parity

•   The absolute purchasing power parity theory (APPPT) predicts that price levels will be the same across countries.
•   Recall that the law of one price states that the same products will have the same prices everywhere.
•   The APPPT applies the same logic to all prices in the country. Basically, it predicts that the cost of living across countries should be the same.

•   It is clear to anyone who has visited a foreign country that the APPPT does not hold.

•   One popular macroeconomic analysis metric to compare economic productivity and standards of living between countries is purchasing power parity (PPP).

•   PPP is an economic theory that compares different countries' currencies through a "basket of goods" approach.

•   According to this concept, two currencies are in equilibrium—known as the currencies being at par—when a basket of goods is priced the same in both countries, taking into account the exchange rates.

Calculating Purchasing Power Parity

•   S=P2/P1

•   where:
•   S= Exchange rate of currency 1 to currency 2
•   P1= Cost of good X in currency 1
•   P2= Cost of good X in currency 2

•   P1= 15%
= 15/100
= 0.15
•   P2= 10%
=10/100
=0.10
S=p2/p1
S=0.15/0.10
S=1.5


3.   Asset Approach by considering an equilibrium in Foreign Exchange Rate Market

The interest parity condition can be used to develop a model of exchange rate determination.
That is, investor behaviour in asset markets which generates interest parity can also explain why the exchange rate may rise and fall in response to market changes.

The first step is to reinterpret the rate of return calculation described above in more general (aggregate) terms.
Thus instead of using the interest rate on a one year CD, we will interpret the interest rates in the two countries as the average interest rates currently prevailing. Similarly, we will imagine that the expected exchange rate is the average expectation across many different individual investors.
The rates of return then are the average expected rates of return on a wide variety of assets between the two countries.
Consider the market for Country 1 in Country 2 depicted in the adjoining diagram. We measure the supply and demand of Country 1s along the horizontal axis and the price of Country 1s (i.e. the exchange rate E Country 1/ Country 2) on the vertical axis. Let S Country 1 represent the supply of Country 1s in exchange for dollars at all different exchange rates that might prevail. The supply is generally by British investors who demand dollars to purchase dollar denominated assets. However, supply of Country 1s might also come from Country 2 investors who decide to convert previously acquired Country 1 currency. Let D Country 1 the demand for Country 1s in exchange for Country 2 at all different exchange rates that might prevail. The demand is generally by Country 2 investors who supply Country 2 to purchase Country 1 denominated assets. Of course, demand might also come from Country 1 investors who decide to convert previously purchased Country 2.

RoR1 = Country 1 + (1 + Country 1) E1 – E2/ E12

E1 – Country 1
E2 – Country 2ÉS country, E --- Country 1) & Country 2)